# Order-2 Delaunay Triangulations Optimize Angles

### v3

@@ 4.3. Global Optimality
The genericity of $A$ implies $\measuredangle xay\neq\measuredangle xby$ whenever points $a,b\in A$ lie on the same side of the line that passes through $x,y\in A$. However, there may be two or more triplets of points in $A$ that define the same angle. It will be convenient to have distinct angles, so we first apply a perturbation that preserves the order of unequal angles while making equal angles different. The relation for the perturbed points implies the same but possibly non-strict relation for the original points, since undoing the perturbation does not change the order of any two angles. So assume that the angles defined by the points in $A$ are distinct, and to derive a contradiction, assume ${{\rm Vector}({D})}\prec{{\rm Vector}({P})}$. More specifically, we write $\alpha_{1}<\alpha_{2}<\ldots<\alpha_{3m}$ and $\varphi_{1}<\varphi_{2}<\ldots<\varphi_{3m}$ for the angles of $D$ and $P$, respectively, and we assume $\alpha_{i}=\varphi_{i}$, for $11\leq i\leq p-$ 1, and $\alpha_{p}<\varphi_{p}$, for some $1\leq p\leq 3m$. In other words, $p$ is the first index at which the two angle vectors differ, and the $p$-th angle of $D$ is smaller than the $p$-th angle of $P$. Write $\alpha=\alpha_{p}$ and let $bac\in D$ be the triangle with $\alpha=\measuredangle bac$. By the assumption of distinct angles, $bac\not\in P$. To simplify the discussion, we assume that the line, $L$, that passes through $b$ and $c$ is horizontal and that $a$ lies above $L$. Since $P$ is a complete triangulation of $A$, every point of $A$ is a vertex of $P$. We consider two cases for the triangles of $P$ that cover the upper side of $bc$.
@@
If there are at least two triangles of $P$ covering the upper side of $bc$, then we consider the triangle $bxy$ that shares $b$ with $bac$. Here $x$ is above $L$ and $y$ is below $L$. Since $c$ is a vertex of $P$, $xy$ intersects $bc$ and $\measuredangle bxy<\measuredangle bxc<\alpha$. As in the first case, this implies that $bxy$ is a triangle in $D$, which leads to a contradiction as triangles $bxy$ and $bac$ intersect. ∎
@@
Let $A\subseteq{{\mathbb{R}}}^{2}$ be finite and generic, $P$ a complete triangulation of $A$, $\Phi_{2}=f(P)$, and $\Delta_{2}=f({{\rm Del}{({A})}})$. Then ${{\rm Vector}({\Phi_{2}})}\preceq{{\rm Vector}({\Delta_{2}})}$.
@@
see Figures 3 and 4. We first consider the case in which $bac$ is a black triangle. There are three subcases, and in each we get a contradiction by constructing two triangles that share interior points. Note that two white triangles may share interior points, but not if they triangulate the same star.
@@
Case 1.2.2: both $x$ and $x^{\prime}$ are equal to $a$. Then $bay,cay^{\prime}\in{{\rm Black}({\Phi_{2}})}$. Since $\measuredangle bay<\alpha$ and $\measuredangle cay^{\prime}<\alpha$, both are white triangles in $\Delta_{2}$. By Lemma 4.3 (1), $bay\in{{\rm White}({\Delta_{2},c})}$ and $cay^{\prime}\in{{\rm White}({\Delta_{2},b})}$, which implies that $cy$ and $by^{\prime}$ are edges in ${{\rm Del}{({A})}}$. If $y\neq y^{\prime}$, then there are three possible choices for the points $b$, $c$, $y$, $y^{\prime}$. First, they form a convex quadrangle, $byy^{\prime}c$, with the points ordered as they are seen from $a$. But then $by^{\prime}$ and $cy$ cross, which contradicts that they both belong to ${{\rm Del}{({A})}}$. Second, $y$ lies inside $bcy^{\prime}$. Since $cay^{\prime}\in{{\rm White}({\Delta_{2},b})}$, the circumcircle of $cay^{\prime}$ encloses $b$ and therefore $y$, which is one point too many for a white triangle in $\Delta_{2}$. Third, $y^{\prime}$ lies inside $bcy$, but this is symmetric to the second choice. Since we get a contradiction for all three choices, we conclude that $y=y^{\prime}$.
@@
Case 2.1: $d$ lies below $L$; see the left and middle panels of Figure 4. Then $\measuredangle bxc<\measuredangle bac$ for all $x\in A$ above $L$, and $\measuredangle byc<\measuredangle bdc$ for all $y\in A$ below $L$. Similar to Case 1.1, we distinguish between the upper side of $bc$ being covered by one or requiring two or more black triangles in $\Phi_{2}$. In both cases, we derive a contradiction by constructing triangles in ${{\rm White}({\Delta_{2},a})}$ that share interior points.
@@
Since $bd$ and $cd$ are edges of triangles in ${{\rm Black}({\Phi_{2}})}$, we have $y=y^{\prime}=d$. Consider ${{\rm st}({P,d})}$, which contains $b$ and $c$ on its boundary. The black triangles in $\Phi_{2}$ that cover the upper side of $bc$ all share $d$ as a vertex, which implies that $bc$ lies inside this star. Indeed, by Lemma 3.4, it is an edge of a triangle in ${{\rm White}({\Phi_{2},d})}$. Thus, there exists a triangle $bzc\in{{\rm White}({\Phi_{2},d})}$ with $z$ above $L$. We have $z\neq a$ by assumption on $\alpha$, so $\measuredangle bzc<\alpha$, which implies that $bzc$ is also a white triangle in $\Delta_{2}$, and since its circumcircle encloses $a$, $bzc\in{{\rm White}({\Delta_{2},a})}$.
To construct a second such white triangle, note that this implies that $ab$ and $ac$ are edges of triangles in ${{\rm Black}({\Delta_{2}})}$. As illustrated in the middle panel of Figure 4, all of $ab,ac,ad,bd,cd$ are edges of black triangles in $\Delta_{2}$, so $bac,bdc\in{{\rm Black}({\Delta_{2}})}$. Hence, $bd$ and $cd$ are edges in the boundary of ${{\rm st}({D,a})}$, and since $bzc\in{{\rm White}({\Delta_{2},a})}$, we also have $bdc\in{{\rm White}({\Delta_{2},a})}$. The angle at $b$ satisfies $\measuredangle dbc<\measuredangle dac<\alpha$ because $a$ lies inside the circumcircle of $dbc$, and since $dbc$ is a triangle in $\Delta_{2}$, it must therefore also be a triangle in $\Phi_{2}$. It cannot be in ${{\rm Black}({\Phi_{2}})}$ because the upper side of $bc$ requires at least two black triangles of $\Phi_{2}$ to be covered, by assumption. Hence, $dbc$ is white in $\Phi_{2}$. It shares interior points with the two black triangles with common edge $dz$ in $\Phi_{2}$, so $dbc\in{{\rm White}({\Phi_{2},z})}$, by Lemma 4.3 (3).
@@ 4.4. Counterexamples
Order beyond 2. Four points in convex position permit only two triangulations: $D={{\rm Del}{({A})}}$, and $P$, which consists of the other two triangles spanned by the four points. As illustrated in Figure 5, ${{\rm Del}_{2}{({A})}}$ consists of shrunken and possibly inverted copies of all four triangles, and ${{\rm Del}_{3}{({A})}}$ consists of shrunken and inverted copies of the two triangles in $P$. Assuming $A$ is generic, Sibson’s theorem implies ${{\rm Vector}({P})}\prec{{\rm Vector}({D})}$. There are two level-$3$ hypertriangulations: the order-$3$ Delaunay triangulation, with ${{\rm Vector}({{{\rm Del}_{3}{({A})}}})}={{\rm Vector}({P})}$, and another, with ${{\rm Vector}({P_{3}})}={{\rm Vector}({D})}$. Hence, ${{\rm Vector}({{{\rm Del}_{3}{({A})}}})}\prec{{\rm Vector}({P_{3}})}$. In words, the vector inequality asserted in Theorem 4.5 for order-$2$ Delaunay trinagulations does not even extend to order $3$.
@@ 5.2. Level-2 Hypertriangulations
Let $P$ be the underlying (order-1) triangulation of $A$, which consists of the images of the black triangles in $P_{2}$ under the inverse aging function. We begin by establishing that $P$ is maximal and therefore $P_{2}$ is complete. Suppose $x\in A$ is not a vertex of $P$, and let $abc$ be the triangle in $P$ that contains $x$ in its interior. Consider the triangle with vertices $c^{\prime}={[{ab}]}$, $b^{\prime}={[{ac}]}$, and $a^{\prime}={[{bc}]}$ in ${{\rm Black}({P_{2}})}$. The edge connecting $b^{\prime}$ and $c^{\prime}$ is shared with ${[{{{\rm wh}({P_{2},a})}}]}$, and this white region contains $x^{\prime}={[{ax}]}$. Since $P_{2}$ is maximal, by assumption, $x^{\prime}$ is a vertex of the restriction of $P_{2}$ to this white region. Recall that the triangle $b^{\prime}d^{\prime}c^{\prime}$ in the constrained Delaunay triangulation of the white region has the property that the angle at $d^{\prime}$ is maximal over all possible choices of $d^{\prime}$ visible from $b^{\prime}$ and $c^{\prime}$. Hence, $\measuredangle b^{\prime}d^{\prime}c^{\prime}\geq\measuredangle b^{\prime}x^{\prime}c^{\prime}$, but also $\measuredangle b^{\prime}x^{\prime}c^{\prime}=\measuredangle bxc>\measuredangle bac=\measuredangle b^{\prime}a^{\prime}c^{\prime}$ because $x$ is inside $abc$. This implies $\measuredangle b^{\prime}d^{\prime}c^{\prime}>\measuredangle b^{\prime}a^{\prime}c^{\prime}$, which contradicts (bw) for $P_{2}$, so $P$ is necessarily maximal.
Applying Lemma 5.3 to $P_{2}$, we conclude that $P$ satisfies (ww). Since $P$ is a maximal, the only choice left is that $P$ is the Delaunay triangulation of $A$. The black triangles in $P_{2}$ thus coincide with the black triangles in the order-$2$ Delaunay triangulation of $A$, and $P_{2}$ restricted to each of its white regions is the constrained Delaunay triangulation of this region. Hence, $P_{2}$ is the order-$2$ Delaunay triangulation of $A$. ∎
@@ 6. Concluding Remarks
In this last section, we discuss open questions about hypertriangulations. The obvious one is whether optimality properties other than angles can be generalized from level $1$ to higher levels: for example the smallest circumcircle [4], the smallest enclosing circle [18], roughness [19], and other functionals [6, Chapter 3] and [15], which are all optimized by the order-$1$ Delaunay triangulation. In addition, we list a small number of more specific questions and conjectures directly related to the discussions in the technical sections of this paper.
@@
Maximal and maximum hypertriangulations. Recall that a hypertriangulation is maximal if no other hypertriangulation of the same level subdivides it. We say a hypertriangulation is maximum if no other hypertriangulation of the same level has more triangles. In an attempt to generalize Lemma 2.6 to levels beyond $2$, we conjecture that the number of triangles in a maximum hypertriangulation depends on the given points but not on how these points are triangulated.

### v4

@@ 4.3. Global Optimality
The genericity of $A$ implies $\measuredangle xay\neq\measuredangle xby$ whenever points $a,b\in A$ lie on the same side of the line that passes through $x,y\in A$. However, there may be two or more triplets of points in $A$ that define the same angle. It will be convenient to have distinct angles, so we first apply a perturbation that preserves the order of unequal angles while making equal angles different. The relation for the perturbed points implies the same but possibly non-strict relation for the original points, since undoing the perturbation does not change the order of any two angles. So assume that the angles defined by the points in $A$ are distinct, and to derive a contradiction, assume ${{\rm Vector}({D})}\prec{{\rm Vector}({P})}$. More specifically, we write $\alpha_{1}<\alpha_{2}<\ldots<\alpha_{3m}$ and $\varphi_{1}<\varphi_{2}<\ldots<\varphi_{3m}$ for the angles of $D$ and $P$, respectively, and we assume $\alpha_{i}=\varphi_{i}$, for $11\leq i\leq p-$ 1, and $\alpha_{p}<\varphi_{p}$, for some $1\leq p\leq 3m$. In other words, $p$ is the first index at which the two angle vectors differ, and the $p$-th angle of $D$ is smaller than the $p$-th angle of $P$. Write $\alpha=\alpha_{p}$ and let $bac\in D$ be the triangle with $\alpha=\measuredangle bac$. By the assumption of distinct angles, $bac\not\in P$. To simplify the discussion, we assume that the line, $L$, that passes through $b$ and $c$ is horizontal and that $a$ lies above $L$. Since $P$ is a complete triangulation of $A$, every point of $A$ is a vertex of $P$. We consider two cases for the triangles of $P$ that cover the upper side of $bc$ [^1] We use the phrase “to cover the upper side of $bc$” informally, in the sense that we consider triangles that intersect or contain $bc$ and some parts of an open neighborhood of $bc$ above $bc$. A more formal explanation is given in the proof of Theorem 4.5..
@@
If there are at least two triangles of $P$ covering the upper side of $bc$, then we consider the triangle $bxy$ that shares $b$ with $bac$. Here $x$ is above $L$ and $y$ is below $L$. Since $c$ is a vertex of $P$, $xy$ intersects $bc$ and $\measuredangle bxy<\measuredangle bxc<\alpha$. As in the first case, this implies that $bxy$ is a triangle in $D$, which leads to a contradiction as the interiors of the triangles $bxy$ and $bac$ overlap. ∎
@@
Let $A\subseteq{{\mathbb{R}}}^{2}$ be finite and generic, $P_{2}$ any complete level-2 hypertriangulation of $A$, and $\Delta_{2}=f({{\rm Del}{({A})}})$. Then ${{\rm Vector}({P_{2}})}\preceq{{\rm Vector}({\Delta_{2}})}$.
@@
Let $P$ be the complete triangulation of $A$ such that $P_{2}=F(P)$. Note, however, that we do not assume that the white triangles of $P_{2}$ define constrained Delaunay triangulations of the white regions. Let $\Phi_{2}=F(P)$ be the designated aging of $P$ so $\Phi_{2}$ shares its black triangles with $P_{2}$, and the white triangles of $\Phi_{2}$ form constrained Delaunay triangulation of the white regions. For each white region of $P_{2}$ and $\Phi_{2}$, we apply Lawson’s edge-flip algorithm to turn the white triangles of $P_{2}$ into the white triangles of $\Phi_{2}$; see [3, Section 2.10.3] for more details about this algorithm. Since every flip lexicographically improves the angle vector, we get ${{\rm Vector}({P_{2}})}\preceq{{\rm Vector}({\Phi_{2}})}$, so it remains to show that ${{\rm Vector}({\Phi_{2}})}\preceq{{\rm Vector}({\Delta_{2}})}$.
@@
see Figures 4 and 5. We first consider the case in which $bac$ is a black triangle. There are three subcases, and in each we get a contradiction by constructing two triangles that share interior points. Note that two white triangles may share interior points, but not if they triangulate the same star.
@@
Case 1.2.2: both $x$ and $x^{\prime}$ are equal to $a$. Then $bay,cay^{\prime}\in{{\rm Black}({\Phi_{2}})}$. Since $\measuredangle bay<\alpha$ and $\measuredangle cay^{\prime}<\alpha$, both are white triangles in $\Delta_{2}$. By Lemma 4.3 (1), $bay\in{{\rm White}({\Delta_{2},c})}$ and $cay^{\prime}\in{{\rm White}({\Delta_{2},b})}$, which implies that $cy$ and $by^{\prime}$ are edges in ${{\rm Del}{({A})}}$. If $y\neq y^{\prime}$, then there are three possible choices for the points $b$, $c$, $y$, $y^{\prime}$. First, they form a convex quadrangle, $byy^{\prime}c$, with the points ordered as they are seen from $a$. The vertices of the quadrangle must be ordered that way as both triangles, $bay$ and $cay^{\prime}$, cover parts of the upper side of $bc$ and cannot overlap since they both belong to ${{\rm Black}({\Phi_{2}})}$. But then $by^{\prime}$ and $cy$ cross, which contradicts that they both belong to ${{\rm Del}{({A})}}$. Second, $y$ lies inside $bcy^{\prime}$. Since $cay^{\prime}\in{{\rm White}({\Delta_{2},b})}$, the circumcircle of $cay^{\prime}$ encloses $b$ and therefore $y$, which is one point too many for a white triangle in $\Delta_{2}$. Third, $y^{\prime}$ lies inside $bcy$, but this is symmetric to the second choice. Since we get a contradiction for all three choices, we conclude that $y=y^{\prime}$.
@@
Case 2.1: $d$ lies below $L$; see the left and middle panels of Figure 5. Then $\measuredangle bxc<\measuredangle bac$ for all $x\in A$ above $L$, and $\measuredangle byc<\measuredangle bdc$ for all $y\in A$ below $L$. Indeed, since $bac\in{{\rm White}({\Delta_{2},d})}$, the circumcircle of $bac$ contains only $d$ inside, because only such white triangles are included in the order-2 Delaunay triangulation $\Delta_{2}$; see Definition 3.1. This implies the inequality for every $x\in A$ above $bc$. For every $y\in A$ below $bc$, we have $\measuredangle bac+\measuredangle byc<\pi$, since $y$ is outside the circumcircle $abc$, while $\measuredangle bac+\measuredangle bdc>\pi$, since $d$ is inside the circumcircle $abc$. This implies $\measuredangle byc<\measuredangle bdc$.
Similar to Case 1.1, we distinguish between the upper side of $bc$ being covered by one or requiring two or more black triangles in $\Phi_{2}$. In both cases, we derive a contradiction by constructing triangles in ${{\rm White}({\Delta_{2},a})}$ that share interior points.
@@
Since $bd$ and $cd$ are edges of triangles in ${{\rm Black}({\Phi_{2}})}$, we have $y=y^{\prime}=d$. Consider ${{\rm st}({P,d})}$, which contains $b$ and $c$ on its boundary. The black triangles in $\Phi_{2}$ that cover the upper side of $bc$ all share $d$ as a vertex, which implies that $bc$ lies inside this star. Indeed, by Lemma 3.4, it is an edge of a triangle in ${{\rm White}({\Phi_{2},d})}$. Indeed, among the points of $A$, the circumcircle of $bac$ contains only $d$ inside, but $d$ is not a vertex of the polygonal region ${{\rm wh}({\Phi_{2},d})}$. Thus, there exists a triangle $bzc\in{{\rm White}({\Phi_{2},d})}$ with $z$ above $L$. We have $z\neq a$ by assumption on $\alpha$, so $\measuredangle bzc<\alpha$, which implies that $bzc$ is also a white triangle in $\Delta_{2}$, and since its circumcircle encloses $a$, $bzc\in{{\rm White}({\Delta_{2},a})}$.
To construct a second such white triangle, note that this implies that $ab$ and $ac$ are edges of triangles in ${{\rm Black}({\Delta_{2}})}$. As illustrated in the middle panel of Figure 5, all of $ab,ac,ad,bd,cd$ are edges of black triangles in $\Delta_{2}$, so $bad,cad\in{{\rm Black}({\Delta_{2}})}$. Hence, $bd$ and $cd$ are edges in the boundary of ${{\rm st}({D,a})}$, and since $bzc\in{{\rm White}({\Delta_{2},a})}$, we also have $bdc\in{{\rm White}({\Delta_{2},a})}$. The angle at $b$ satisfies $\measuredangle dbc<\measuredangle dac<\alpha$ because $a$ lies inside the circumcircle of $dbc$, and since $dbc$ is a triangle in $\Delta_{2}$, it must therefore also be a triangle in $\Phi_{2}$. It cannot be in ${{\rm Black}({\Phi_{2}})}$ because the upper side of $bc$ requires at least two black triangles of $\Phi_{2}$ to be covered, by assumption. Hence, $dbc$ is white in $\Phi_{2}$. It shares interior points with the two black triangles with common edge $dz$ in $\Phi_{2}$, so $dbc\in{{\rm White}({\Phi_{2},z})}$, by Lemma 4.3 (3).
@@ 4.4. Counterexamples
Order beyond 2. Four points in convex position permit only two triangulations: $D={{\rm Del}{({A})}}$, and $P$, which consists of the other two triangles spanned by the four points. As illustrated in Figure 6, ${{\rm Del}_{2}{({A})}}$ consists of shrunken and possibly inverted copies of all four triangles, and ${{\rm Del}_{3}{({A})}}$ consists of shrunken and inverted copies of the two triangles in $P$. Assuming $A$ is generic, Sibson’s theorem implies ${{\rm Vector}({P})}\prec{{\rm Vector}({D})}$. There are two level-$3$ hypertriangulations: the order-$3$ Delaunay triangulation, with ${{\rm Vector}({{{\rm Del}_{3}{({A})}}})}={{\rm Vector}({P})}$, and another, with ${{\rm Vector}({P_{3}})}={{\rm Vector}({D})}$. Hence, ${{\rm Vector}({{{\rm Del}_{3}{({A})}}})}\prec{{\rm Vector}({P_{3}})}$. In words, the vector inequality asserted in Theorem 4.5 for order-$2$ Delaunay triangulations does not even extend to order $3$.
@@ 5.2. Level-2 Hypertriangulations
Let $P$ be the underlying (order-1) triangulation of $A$, which consists of the images of the black triangles in $P_{2}$ under the inverse aging function. We begin by establishing that $P$ is maximal and therefore $P_{2}$ is complete. Suppose $x\in A$ is not a vertex of $P$, and let $abc$ be the triangle in $P$ that contains $x$ in its interior. Consider the triangle with vertices $c^{\prime}={[{ab}]}$, $b^{\prime}={[{ac}]}$, and $a^{\prime}={[{bc}]}$ in ${{\rm Black}({P_{2}})}$. The edge connecting $b^{\prime}$ and $c^{\prime}$ is shared with ${[{{{\rm wh}({P_{2},a})}}]}$, and this white region contains $x^{\prime}={[{ax}]}$. Since $P_{2}$ is maximal, by assumption, $x^{\prime}$ is a vertex of the restriction of $P_{2}$ to this white region. Let $d^{\prime}$ be the vertex of $P_{2}$ such that $b^{\prime}d^{\prime}c^{\prime}$ is a triangle of ${{\rm White}({P_{2},a})}$. This triangle belongs to the constrained Delaunay triangulation of the white region, because this is the only triangulation that satisfies the (ww) property for all interior edges. Recall that the triangle $b^{\prime}yc^{\prime}$ in the constrained Delaunay triangulation has the property that the angle at $y$ is maximal over all possible choices of $y\in A$ in the region visible from $b^{\prime}$ and $c^{\prime}$. Hence, $\measuredangle b^{\prime}d^{\prime}c^{\prime}\geq\measuredangle b^{\prime}x^{\prime}c^{\prime}$, as $x^{\prime}$ is visible from $b^{\prime}$ and $c^{\prime}$, because the whole triangle $b^{\prime}x^{\prime}c^{\prime}$ lies in the white region, but also $\measuredangle b^{\prime}x^{\prime}c^{\prime}=\measuredangle bxc>\measuredangle bac=\measuredangle b^{\prime}a^{\prime}c^{\prime}$, because $x$ is inside $abc$. This implies $\measuredangle b^{\prime}d^{\prime}c^{\prime}>\measuredangle b^{\prime}a^{\prime}c^{\prime}$, which contradicts (bw) for $P_{2}$, so $P$ is necessarily maximal.
Applying Lemma 5.3 to $P_{2}$, we conclude that $P$ satisfies (ww). Since $P$ is maximal, the only choice left is that $P$ is the Delaunay triangulation of $A$. The black triangles in $P_{2}$ thus coincide with the black triangles in the order-$2$ Delaunay triangulation of $A$, and $P_{2}$ restricted to each of its white regions is the constrained Delaunay triangulation of this region. Hence, $P_{2}$ is the order-$2$ Delaunay triangulation of $A$. ∎
@@ 6. Concluding Remarks
In this last section, we discuss open questions about hypertriangulations. The obvious one is whether optimality properties other than angles can be generalized from level $1$ to higher levels: for example the smallest circumcircle [5], the smallest enclosing circle [19], roughness [20], and other functionals [7, Chapter 3] and [16], which are all optimized by the order-$1$ Delaunay triangulation. In addition, we list a small number of more specific questions and conjectures directly related to the discussions in the technical sections of this paper.
@@
Maximal and maximum hypertriangulations. Recall that a hypertriangulation is maximal if no other hypertriangulation of the same level subdivides it. We say a hypertriangulation is maximum if no other hypertriangulation of the same level has more triangles. In an attempt to generalize Lemma 2.6 to levels beyond $2$, we conjecture that the number of triangles in a maximal hypertriangulation depends on the given points but not on how these points are triangulated.